# Targeted collapse regularized autoencoder for anomaly detection: black hole at the center

### v1

@@ Abstract
Autoencoders have been extensively used in the development of recent anomaly detection techniques. The premise of their application is based on the notion that after training the autoencoder on normal training data, anomalous inputs will exhibit a significant reconstruction error. Consequently, this enables a clear differentiation between normal and anomalous samples. In practice, however, it is observed that autoencoders can generalize beyond the normal class and achieve a small reconstruction error on some of the anomalous samples. To improve the performance, various techniques propose additional components and more sophisticated training procedures. In this work, we propose a remarkably straightforward alternative: instead of adding neural network components, involved computations, and cumbersome training, we complement the reconstruction loss with a computationally light term that regulates the norm of representations in the latent space. The simplicity of our approach minimizes the requirement for hyperparameter tuning and customization for new applications which, paired with its permissive data modality constraint, enhances the potential for successful adoption across a broad range of applications. We test the method on various visual and tabular benchmarks and demonstrate that the technique matches and frequently outperforms alternatives. We also provide a theoretical analysis and numerical simulations that help demonstrate the underlying process that unfolds during training and how it can help with anomaly detection. This mitigates the black-box nature of autoencoder-based anomaly detection algorithms and offers an avenue for further investigation of advantages, fail cases, and potential new directions.
@@ 1 Introduction
Anomaly detection is the task of identifying instances of data that significantly deviate from normal [^1] The term “normal” in this text indicates “not anomalous.” observations. The task has many important applications including fault detection purarjomandlangrudi2014data ; yan2023hybrid ; liang2020sparse ; zhang2022anomaly ; fan2020defective ; chen2023adaptive , medical diagnosis fernando2021deep ; han2021madgan ; kascenas2022denoising ; tschuchnig2022anomaly , fraud detection pourhabibi2020fraud ; pang2019deep ; pang2019deviation ; zheng2019one ; hilal2022financial , performance and throughput anomaly detection and bottleneck identificationibidunmoye2015performance ; cherkasova2009automated ; chen2023hierarchical , and network intrusion detection ahmed2016survey ; al2018deep ; dutta2020deep ; aygun2017network ; wang2020cloud ; lokman2019deep ; naseer2018enhanced ; yang2020real , among others. Considering that anomalous samples are rare and can often highly vary, anomaly detection is typically done through learning regularities present in normal data. Subsequently, samples that reside far outside the normal class representation are regarded as anomalous.
@@
Deep autoencoders hinton2006reducing ; bengio2006greedy are powerful tools to learn unsupervised representations of high dimensional data and have been central to many algorithms among various modern techniques for anomaly detection sakurada2014anomaly ; zhou2017anomaly ; gong2019memorizing ; gao2022tsmae . The premise for using autoencoders in this context is that when trained on normal samples, they fail to effectively reconstruct anomalous samples as by definition they lie outside the training distribution. It therefore is possible to devise an anomaly scoring based on the reconstruction error. This premise, however, does not always hold and there indeed can be cases where an anomalous input can be well reconstructed perera2019ocgan ; zong2018deep .
@@
Various strategies have been proposed in the literature to improve the performance of autoencoders by including additional components and more sophisticated training procedures. Some examples of this include adding a memory module gong2019memorizing ; gao2022tsmae ; yang2021memory , visual and latent discriminators yang2021memory ; perera2019ocgan , classifier perera2019ocgan , adversarial training yang2021memory ; perera2019ocgan , and negative mining perera2019ocgan . The added modules and involved training procedures can lead to increased costs associated with the utilization of such models, potentially making them less accessible to various applications and more challenging to troubleshoot.
@@
In this work, we adopt two principles and aim to shed light on the mechanisms by which they enhance the utility of autoencoders in anomaly detection. First, we promote feature compactness in the latent space, ensuring that normal samples exhibit similar representations; secondly, we require feature descriptiveness so that the latent representation retains relevant information for effective separation between normal and anomalous samples. Based on these principles, we propose an autoencoder regularization that penalizes the norm of bottleneck representation. This is a simple and computationally efficient modification especially given the lower dimensionality of the latent space. We will show that this simple adjustment makes the anomaly detection performance of an encoder-decoder architecture match or surpass that of many deep techniques that incorporate the complex modifications mentioned earlier. We also theoretically analyze a simplified case and use the results to offer an explanation regarding the effects of norm minimization on learning dynamics and how it might help with anomaly detection.
@@
The rest of the paper is organized as follows: in Section 2 we discuss related work on anomaly detection and autoencoder regularization. Section 3 discusses the proposed methodology and theoretical analysis, as well as empirical demonstrations for a simple case. In Section 4 we present anomaly detection results on benchmark datasets. Section 5 concludes the paper.
@@ 2 Related Work
Regularized autoencoders have shown improvements over their unregularized counterparts, resulting in better performance on various applications due to the advantages they introduce alain2014regularized ; guo2016deep ; makhzani2013k ; vincent2010stacked ; rifai2011contractive ; kingma2013auto . Sparse autoencoders makhzani2013k regularize the network to encourage sparsity in hidden layers, which serves as an information bottleneck. Denoising autoencoders vincent2010stacked promote robustness to corrupted data and better expressive features by training the model to reconstruct a denoised version of the noisy input. Expanding on a similar intuition, contractive autoencoders rifai2011contractive promote feature robustness to variations in the neighborhood of training instances by penalizing the norm of the Jacobian matrix of the encoder’s activations with respect to input samples. Variational autoencoders kingma2013auto regularize the training by introducing the prior that additionaly incentivizes samples to follow a standard Gaussian distribution in latent space and thus learns an approximation of data distribution parameters rather than arbitrary functions for encoding and reconstruction. These variants of autoencoders have been extensively used in applications of anomaly detection sakurada2014anomaly ; zhou2017anomaly including fault detection and industrial health monitoring yan2023hybrid ; liang2020sparse ; zhang2022anomaly ; fan2020defective , network intrusion, anomaly, and cyber-attack detection al2018deep ; dutta2020deep ; aygun2017network ; wang2020cloud ; lokman2019deep ; naseer2018enhanced ; yang2020real ; nguyen2019gee , video anomaly detection narasimhan2018dynamic , medical diagnosis kascenas2022denoising , geochemical exploration xiong2021robust , and high-energy physics searches finke2021autoencoders .
@@
To improve the performance of autoencoders in anomaly detection, some recent works propose using external memory to learn and record prototypical normal patterns gong2019memorizing ; gao2022tsmae ; yang2021memory . The encoding is then used as a query to read the most relevant items from the memory based on an attention mechanism. The combined memory items are then used for reconstruction. yang2021memory demonstrates that by incorporating bidirectional generative adversarial networks (BiGANs) donahue2016adversarial with adversarial training and adding a cycle-consistency loss, memory units improve and tend to lie on the boundary of the convex hull of normal data encodings.
@@
One-Class GAN (OCGAN) perera2019ocgan focuses on ensuring that anomalous samples are poorly reconstructed by bounding the latent space and ensuring various regions of it will only represent examples of the normal class and only normal class samples can be generated from it through the decoder. To do this, the algorithm bounds the latent space and incorporates adversarial training with latent and visual discriminators, as well as a classifier that is used to ensure generated samples belong to the normal class. In addition, perera2019ocgan leverages negative mining to actively seek and improve upon regions of latent space that produce samples of poor quality.
@@
Besides these works that use adversarial training mainly to aid with the reconstruction-based anomaly score and autoencoder performance, there are many works based on GANs and adversarial training that explicitly use the discriminator output to complement or entirely substitute the reconstruction error schlegl2017unsupervised ; zhao2016energy ; zenati2018adversarially ; akcay2019ganomaly . There are also techniques that attempt to directly learn the anomaly score. This can, for instance, be based on training an ordinal regression neural network that gives different scores to sample pairs given the pair composition pang2019deep . Differently, likelihood-based techniques operate on the premise that normal and anomalous instances respectively correspond to high and low-probability events and therefore the model is optimized to, for instance, obtain high probability for training samples using a softmax or a noise contrastive estimate (NCE) objective gutmann2010noise ; chen2016entity .
@@ 3.1 Problem Statement
where $En(\cdot;\theta_{e})$ and $De(\cdot;\theta_{d})$ are the encoder and decoder, respectively parameterized by $\theta_{e}$ and $\theta_{d}$. It is then expected that normal samples (i.e. samples from the same distribution as training) incur smaller reconstruction errors compared to anomalous, out-of-distribution samples. Prior research perera2019ocgan ; zong2018deep , however, shows that the assumptions on reconstruction quality might be dubious and the effectiveness of anomaly scores can vary depending on the specific type of anomaly. If the anomalous sample shares some critical features with normal samples, it may still be well reconstructed.
@@ 3.2 Targeted Collapse Regularized Autoencoders
The logic behind the adoption of this regularization is rather straightforward: due to the underlying similarity across normal samples, they should reside closely together in the latent space. A similar inductive bias is, for instance, introduced in the design of Deep Support Vector Data Description (DSVDD) ruff2018deep , where the encoder is tasked with learning a mapping of normal data into a hypersphere with center $c$ and radius $R$ with minimal volume, outside of which anomalies are expected to reside. Unlike in that work, we do not impose a domain constraint for the resulting distribution, but rather only push the encodings towards the origin. This addition stipulates that not only is the autoencoder forced to efficiently reconstruct the normal data, but also to map them such that they form a compact latent representation, emphasizing shared characteristics among the normal samples, thereby capturing their prevalent patterns.
@@
Not all mappings leading to small representation norms are useful in this sense. An encoder’s latent space representations can be arbitrarily scaled down uniformly to reduce the norms. As another example, one can simply obtain a trivial function that maps any input to the origin. In fact, ruff2018deep explicitly discusses the phenomena (referred to in that work as “hypersphere collapse” given the constraint on latent representations) and demonstrates how it imposes restrictions on the choice of activation function and inhibits the ability to incorporate the hypersphere center and network biases as trainable parameters and instead offers an empirical strategy to pick a prespecified fixed center.
With gradient descent training, the latter scenario (collapsed representation) is avoided, noting that such a representation is not an optima: a collapsed latent space results in a strictly positive reconstruction error. As such, during training the encoder is taken away from trivial collapse in the parameter space. As for the former case, in the following we will provide a theoretical analysis of the learning dynamics for a simplified case that also shows we do not merely retrieve a scaled autoencoder bottleneck representation. In fact, the learning dynamics analysis and subsequent numerical simulations demonstrate that the regularization nicely complements the plain autoencoder and increases its utility for anomaly detection, which our results in Section 4 also corroborate.
@@ 3.3 Analysis of Learning Dynamics
Let us now consider the scenario in which only the reconstruction error constitutes the loss of a linear autoencoder. It is known (as shown e.g. in plaut2018principal ) that a linear autoencoder with a $d$-dimensional bottleneck and trained using mean squared error loss tends to encode data in the space spanned by the $d$ principal components (i.e. eigenvectors associated with the $d$ largest eigenvalues of covariance), fully discarding low eigenvalue directions. Consequently, it produces a low-rank approximation of the data at reconstruction.
@@ 4.1 Datasets and Implementation Details
We present the results of Toll on three datasets: MNIST lecun2010mnist , CIFAR-10 krizhevsky2009learning , and Arrhythmia dua2017uci . For MNIST and CIFAR-10 we use the same protocol as that used in e.g. ruff2018deep ; perera2019ocgan ; yang2021memory . MNIST and CIFAR-10 are visual datasets and both have 10 classes and have respectively 60000 and 50000 images in their training set and 10000 images in test sets. Figure 3 shows sample images from all classes in MNIST and CIFAR-10. We also set aside 10000 samples from the training set for validation. In both datasets, we perform 10 experiments, each time considering one of the classes as the normal class and all the other classes as anomalous. Training is done on samples from the normal class of the associated experiment. (i.e. approximately 5000 samples for MNIST and 4000 for CIFAR-10).
Figure 3: Samples from MNIST (a) and CIFAR-10 (b) datasets. Rows show different classes.
@@
MNIST dataset was normalized to have a zero mean and a standard deviation of 1 for pixel values. CIFAR-10 images were preprocessed with global contrast normalization using the $L^{1}$-norm and rescaled to $\left[0,1\right]$ with min-max scaling. The encoder and decoder architectures used for CIFAR-10 experiments are similar to encoder and generator architectures in dumoulin2016adversarially ; yang2021memory . The architectures for MNIST and Arrhythmia experiments are also based on yang2021memory . See Appendix B for detailed architectures.
In all experiments, training is done through iterating over 2000 random batches with checkpoints considered at 20 batch intervals. The batch sizes for MNIST, CIFAR-10, and Arrhythmia are respectively 1000, 50, and 100. Adam optimizer with a fixed learning rate is used for training. Learning rates, $\beta$ values, and the bottleneck dimensions are picked based on the validation set performance. We keep these hyperparameters unchanged across the 10 classes in MNIST and CIFAR-10, but we note that test performances further improve, in many instances considerably, if hyperparameters are picked specific to the normal class. The best-performing model on the validation set is selected for subsequent evaluation on the test set. Each experiment is repeated for 10 different random seeds and averaged results are reported.
We comply with the existing literature for evaluation. The metric used to evaluate results on MNIST and CIFAR-10 is area under ROC curve (AUC). For Arrhythmia, we set the score threshold such that we label the expected 15 percent of samples with the highest anomaly scores as anomalous. We then calculate the $F_{1}$ score.

### v2

@@ Abstract
Autoencoders have been extensively used in the development of recent anomaly detection techniques. The premise of their application is based on the notion that after training the autoencoder on normal training data, anomalous inputs will exhibit a significant reconstruction error. Consequently, this enables a clear differentiation between normal and anomalous samples. In practice, however, it is observed that autoencoders can generalize beyond the normal class and achieve a small reconstruction error on some of the anomalous samples. To improve the performance, various techniques propose additional components and more sophisticated training procedures. In this work, we propose a remarkably straightforward alternative: instead of adding neural network components, involved computations, and cumbersome training, we complement the reconstruction loss with a computationally light term that regulates the norm of representations in the latent space. The simplicity of our approach minimizes the requirement for hyperparameter tuning and customization for new applications which, paired with its permissive data modality constraint, enhances the potential for successful adoption across a broad range of applications. We test the method on various visual and tabular benchmarks and demonstrate that the technique matches and frequently outperforms more complex alternatives. We further demonstrate that implementing this idea in the context of state-of-the-art methods can further improve their performance. We also provide a theoretical analysis and numerical simulations that help demonstrate the underlying process that unfolds during training and how it helps with anomaly detection. This mitigates the black-box nature of autoencoder-based anomaly detection algorithms and offers an avenue for further investigation of advantages, fail cases, and potential new directions.
@@ 1 Introduction
Anomaly detection is the task of identifying instances of data that significantly deviate from normal [^1] The term “normal” in this text indicates “not anomalous.” observations. The task has many important applications including fault detection [51, 63, 36, 69, 19, 10], medical diagnosis [20, 27, 32, 58], fraud detection [50, 45, 46, 71, 28], performance and throughput anomaly detection and bottleneck identification[31, 11, 8], and network intrusion detection [1, 3, 18, 5, 60, 38, 43, 64], among others. Considering that anomalous samples are rare and can often highly vary, anomaly detection is typically done through learning regularities present in normal data. Subsequently, samples that reside far outside the normal class representation are regarded as anomalous.
@@
Deep autoencoders [29, 6] are powerful tools to learn unsupervised representations of high dimensional data and have been central to many algorithms among various modern techniques for anomaly detection [54, 72, 24, 22]. The premise for using autoencoders in this context is that when trained on normal samples, they fail to effectively reconstruct anomalous samples as by definition they lie outside the training distribution. It is therefore possible to devise an anomaly scoring based on the reconstruction error. This premise, however, does not always hold and there indeed can be cases where an anomalous input is well reconstructed [48, 73].
@@
Various strategies have been proposed in the literature to improve the performance of autoencoders by including additional components and more sophisticated training procedures. Some examples of this include adding a memory module [24, 22, 66], visual and latent discriminators [66, 48], classifier [48], adversarial training [66, 48], and negative mining [48]. The added modules and involved training procedures can lead to increased costs associated with the utilization of such models, potentially making them less accessible to various applications and more challenging to troubleshoot.
@@
In this work, we adopt two principles and aim to shed light on the mechanisms by which they enhance the utility of autoencoders in anomaly detection. First, we promote feature compactness in the latent space, ensuring that normal samples exhibit similar representations; second, we require feature descriptiveness so that the latent representation retains relevant information for effective separation between normal and anomalous samples. Based on these principles, we propose a regularization that penalizes the norm of latent representations. This is a simple and computationally efficient modification, especially given the lower dimensionality of the latent space. We will show that this simple adjustment makes the anomaly detection performance of an encoder-decoder architecture match or surpass that of many deep techniques that incorporate the complex modifications mentioned earlier. We also theoretically analyze a simplified case and use the results to offer an explanation regarding the effects of norm minimization on learning dynamics and how it might help with anomaly detection. We further demonstrate that by virtue of promoting compact representations, the proposed regularization can enhance the performance of other techniques that rely on the separation of learned normal/anomalous latent representations.
@@
The rest of the paper is organized as follows: in Section 2 we discuss related work on anomaly detection and autoencoder regularization. Section 3 discusses the proposed methodology and theoretical analysis, as well as empirical demonstrations for a simplified case. In Section 4 we present anomaly detection results on benchmark datasets as well as ablation studies. Section 5 concludes the paper.
@@ 2 Related Work
Regularized autoencoders have shown improvements over their unregularized counterparts, resulting in better performance on various applications due to the advantages they introduce [4, 25, 39, 59, 52, 33]. Sparse autoencoders [39] regularize the network to encourage sparsity in hidden layers, which serves as an information bottleneck. Denoising autoencoders [59] promote robustness to corrupted data and better expressive features by training the model to reconstruct a denoised version of the noisy input. Expanding on a similar intuition, contractive autoencoders [52] promote feature robustness to variations in the neighborhood of training instances by penalizing the norm of the Jacobian matrix of the encoder’s activations with respect to input samples. Variational autoencoders [33] regularize the training by introducing the prior that additionaly incentivizes samples to follow a standard Gaussian distribution in latent space and thus learns an approximation of data distribution parameters rather than arbitrary functions for encoding and reconstruction. These variants of autoencoders have been extensively used in applications of anomaly detection [54, 72] including fault detection and industrial health monitoring [63, 36, 69, 19], network intrusion, anomaly, and cyber-attack detection [3, 18, 5, 60, 38, 43, 64, 44], video anomaly detection [42], medical diagnosis [32], geochemical exploration [62], and high-energy physics searches [21].
@@
To improve the performance of autoencoders in anomaly detection, some recent works propose using external memory to learn and record prototypical normal patterns [24, 22, 66]. The encoding is then used as a query to read the most relevant items from the memory based on an attention mechanism. The combined memory items are then used for reconstruction. [66] demonstrates that by incorporating bidirectional generative adversarial networks (BiGANs) [14] with adversarial training and adding a cycle-consistency loss, memory units improve and tend to lie on the boundary of the convex hull of normal data encodings.
@@
One-Class GAN (OCGAN) [48] focuses on ensuring that anomalous samples are poorly reconstructed by bounding the latent space and ensuring various regions of it will only represent examples of the normal class and only normal class samples can be generated from it through the decoder. To do this, the algorithm bounds the latent space and incorporates adversarial training with latent and visual discriminators, as well as a classifier that is used to ensure generated samples belong to the normal class. In addition, [48] leverages negative mining to actively seek and improve upon regions of latent space that produce samples of poor quality.
@@
Besides these works that use adversarial training mainly to aid with the reconstruction-based anomaly score and autoencoder performance, there are many works based on GANs and adversarial training that explicitly use the discriminator output to complement or entirely substitute the reconstruction error [55, 70, 67, 2]. There are also techniques that attempt to directly learn the anomaly score. This can, for instance, be based on training an ordinal regression neural network that gives different scores to sample pairs given the pair composition [45]. Differently, likelihood-based techniques operate on the premise that normal and anomalous instances respectively correspond to high and low-probability events and therefore the model is optimized to, for instance, obtain high probability for training samples using a softmax or a noise contrastive estimate (NCE) objective [26, 9].
@@ 3.1 Problem Statement
where $En(\cdot;\theta_{e})$ and $De(\cdot;\theta_{d})$ are the encoder and decoder, respectively parameterized by $\theta_{e}$ and $\theta_{d}$. It is then expected that normal samples (i.e. samples from the same distribution as training) incur smaller reconstruction error compared to anomalous, out-of-distribution samples. Prior research [48, 73], however, shows that the assumptions on reconstruction quality might be dubious and the effectiveness of anomaly scores can vary depending on the specific type of anomaly. If the anomalous sample shares some critical features with normal samples, it may still be well reconstructed.
@@ 3.2 Targeted Collapse Regularized Autoencoders
The logic behind the adoption of this regularization is rather straightforward: due to the underlying similarity across normal samples, they should reside closely together in the latent space. A similar inductive bias is, for instance, introduced in the design of Deep Support Vector Data Description (DSVDD) [53], where the encoder is tasked with learning a mapping of normal data into a hypersphere with center $c$ and radius $R$ with minimal volume, outside of which anomalies are expected to reside. Unlike in that work, we do not impose a domain constraint for the resulting distribution, but rather only push the encodings towards the origin. This addition stipulates that not only is the autoencoder forced to efficiently reconstruct the normal data, but also to map them such that they form a compact latent representation, emphasizing shared characteristics among the normal samples, thereby capturing their prevalent patterns.
@@
Not all mappings leading to small representation norms are useful in this sense. An encoder’s latent space representations can be arbitrarily scaled down uniformly to reduce the norms. As another example, one can simply obtain a trivial function that maps any input to the origin. In fact, [53] explicitly discusses the phenomena (referred to in that work as “hypersphere collapse” given the constraint on latent representations) and demonstrates how it imposes restrictions on the choice of activation function and inhibits the ability to incorporate the hypersphere center and network biases as trainable parameters and instead offers an empirical strategy to pick a prespecified fixed center.
By enforcing reconstruction, the latter scenario (collapsed representation) is avoided, noting that such a representation is not an optima: a collapsed latent space results in a strictly positive reconstruction error. As such, during training the encoder is taken away from trivial collapse in the parameter space. As for the former case, in the following we will provide a theoretical analysis of the learning dynamics for a simplified case that also shows we do not merely retrieve a scaled autoencoder bottleneck representation. In fact, the learning dynamics analysis and subsequent numerical simulations demonstrate that the proposed regularization nicely complements the plain autoencoder and increases its utility for anomaly detection, which our results in Section 4 also corroborate.
@@ 3.3 Analysis of Learning Dynamics
Let us now consider the scenario in which only the reconstruction error constitutes the loss of a linear autoencoder. It is known (as shown e.g. in [49]) that a linear autoencoder with a $d$-dimensional bottleneck and trained using mean squared error loss encodes data in the space spanned by the $d$ principal components (i.e. eigenvectors associated with the $d$ largest eigenvalues of covariance), fully discarding low eigenvalue directions. Consequently, it produces a low-rank approximation of the data at reconstruction.
@@ 4.1 Datasets and Implementation Details
We present the results of Toll on five datasets: MNIST [35], Fashion-MNIST [61], CIFAR-10 [34], CIFAR-100 [34], and Arrhythmia [16]. For MNIST, Fashion-MNIST, CIFAR-10, and CIFAR-100 we use the same protocol as that used in e.g. [53, 48, 66]. MNIST, Fashion-MNIST, CIFAR-10, and CIFAR-100 are visual datasets. MNIST, Fashion-MNIST, and CIFAR-10 all have 10 classes. CIFAR-100 consists of 20 superclasses used in this task, each containing 5 closely related classes. For instance, the superclass aquatic mammals contains classes beaver, dolphin, otter, seal, and whale. These datasets respectively have 60000, 60000, 50000, and 50000 images in their training set and 10000 images in test sets. Figure 3 shows sample images from all classes in the four visual datasets. We also set aside 10000 samples from the training set for validation. In all datasets, we perform 10 experiments, each time considering one of the classes as the normal class and all the other classes as anomalous. Training is done on samples from the normal class of the associated experiment. (i.e. approximately 5000 samples for MNIST and Fashion-MNIST, 4000 for CIFAR-10, and 2000 for CIFAR-100).
Figure 3: Samples from MNIST (a, top left), Fashion-MNIST (a, bottom left), CIFAR-100 (a, right), and CIFAR-10 (b) datasets. Rows show different classes.
@@
MNIST and Fashion-MNIST datasets were normalized to have a zero mean and a standard deviation of 1 for pixel values. CIFAR-10 images were preprocessed with global contrast normalization using the $L^{1}$-norm and rescaled to $\left[0,1\right]$ with min-max scaling. The encoder and decoder architectures used for CIFAR-10 and CIFAR-100 experiments are similar to encoder and generator architectures in [17, 66]. The architectures for MNIST, Fashion-MNIST, and Arrhythmia experiments are also based on [66]. See Appendix B for detailed architectures.
In all experiments, training is done through iterating over 2000 random batches with checkpoints considered at 20 batch intervals. The batch sizes for MNIST, Fashion-MNIST, CIFAR-10, CIFAR-100, and Arrhythmia are respectively 1000, 500, 50, 100, and 100. Adam optimizer with a fixed learning rate is used for training. Learning rates, $\beta$ values, and the bottleneck dimensions are picked based on the validation set performance. We keep these hyperparameters unchanged across all normal classes in the four visual datasets, but we note that test performances further improve, in many instances considerably, if hyperparameters are picked specific to the normal class. The best-performing model on the validation set is selected for subsequent evaluation on the test set. Each experiment is repeated for 10 different random seeds and averaged results are reported.
We comply with the existing literature for evaluation. The metric used to evaluate results on MNIST, Fashion-MNIST, CIFAR-10, and CIFAR-100 is the area under ROC curve (AUC). For Arrhythmia, we set the score threshold such that we label the expected 15 percent of samples with the highest anomaly scores as anomalous. We then calculate the $F_{1}$ score.
To further demonstrate the utility of targeted collapse, we also apply the collapse idea in the context of a well-performing anomaly detection method [41]. In this technique, which we refer to as FITYMI (Fake It Till You Make It) following the paper title, the authors first generate synthetic anomalous data using a diffusion model [57] and then fine-tune a binary classifier that incorporates a strong pre-trained backbone. The backbone mainly considered in that work is a pre-trained vision transformer [15]. Specifically, ViT-B/16 pre-trained on ImageNet-21K [13] is adopted.
We replicate the FITYMI experiments on CIFAR-10 and CIFAR-100 datasets in the exact manner described in the paper using the code and generated anomalous samples provided by the authors with no modifications. To then test the targeted collapse idea, we simply add the collapse term for normal samples to the classification loss during training without any other modifications of the parameters and the pipeline. Given the pre-trained backbone in this case, the target for collapse in the latent space was set as the mean of the normal samples, as approximated per batch, in order to simplify the implementation. We again repeat each experiment 10 times and report averaged results.